(b) Mr. lim   is an international investor, just sold a share of UMW a United Kingdom firm for GBP5080. He bought the share for GBP4850 a year ago. The exchange rate is gbp1.50/$ now and was GBP1.48/$ a year ago.Mr.lim received GBP120 as a cash dividend immediately before the share was sold. Calculate the rate on this investment in US Dollar.
Answer #1

Dollar used for purchase of one share = 4850*1.48 = $7,178

Sale + Dividend proceeds in Dollars = (5080+120)*1.50 = $7,800

Rate of return = (Sales proceeds + Dividend - Purchase price)/Purchase price

= (7,800 - 7178)/7178

= 8.6654%

i.e. 8.67%